OpenStudy (anonymous):

A car moving with a velocity of 20 meters/second has 1.8 * 105 joules of kinetic energy. What is the mass of the car

OpenStudy (anonymous):

wat is formuila for kinetic energy of a body?

OpenStudy (anonymous):

KE = (1/2) M V^2. M = 2*KE/V^2. M = 2*1.8*105/20^2 J/(m^2/s^2) = 2*1.8*1.05/400 kg = 0.00945 kg.